OpenStudy (anonymous):

what is the length of a segment with an endpoint of (-5,1) and a midpoint of (-2,3)

OpenStudy (anonymous):

you have a choice find the other endpoint and compute the distance or what is easier is find the distance between these two points and double it

OpenStudy (abb0t):

What is the length of a segment with an endpoint of (-5,1) and a midpoint of (-2,3) horizontal distance from endpoint to midpoint = [(-2) - (-5)] = (3) Vertical distance from end point to midpoint = {3 - 1} = (2) Distance between endpoint and midpoint = \(\sqrt{ 3^2 + 2^2}\) = \(\sqrt{13}\)